Evaluate x² - 4(x – y), given x = 8 and y = 3
Answer by praseenakos@yahoo.com(507)


x² - 4(x – y), given x = 8 and y = 3
8^2 -4(8-3)
=64-4(5)
=64-20
=44 therefore X^2 -4(X-Y) = 44